i am asked to solve using the square root property
(x+6)^2 = 121

Answer by checkley77(12844)

(X+6)^2=121
(X+6)^2=11^2
X+6=11
X=11-6
X=5 ANSWER.
PROOF:
(5+6)^2=121
11^2=121
121=121